# On the stability of global forecasting models

### v1

@@ 3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir & Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Stability metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
@@
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. Following Zanotti (2025b), Table 2 outlines the key parameters used in our experiments, including a test set spanning at least one full year to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency.
Table 2: The retraining scenarios, the test window, and the horizon for the M5 and the VN1 datasets.
Dataset Frequency Retraining Scenarios (r) Test Window (T) Horizon (h)
M5 Daily (7) 7, 14, 21, 30, 60, 90, 120, 150, 180, 364 364 28
VN1 Weekly (52) 1, 2, 3, 4, 6, 8, 10, 13, 26, 52 52 13
Furthermore, to investigate the impact of retraining frequency on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b). Each window $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. Table 2 summarizes the selected retraining windows, test periods, and forecast horizons. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on aligned datasets, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. Following Zanotti (2025b), we treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
@@ 3.4.1 Stability metrics
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and $n-1$ respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply less unstable predictions.
@@
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Conformal Inference is a flexible and robust framework that provides distribution-free, model-agnostic prediction intervals with guaranteed finite-sample coverage under mild assumptions (Vovk et al., 2005). Although initially developed for i.i.d. data, recent extensions have adapted it for time series applications by relaxing the exchangeability assumption (Stankeviciute et al., 2021). Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying levels of retraining frequency. Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic forecasts, we proposed a new measure, the Multi-Quantile Change (MQC, or Multi-Quantile Loss Change, MQLC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
$\text{QC}=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\left(q\cdot(\hat{y}_{t,n-1}-\hat{y}_{t,n})\cdot\mathbb{I}_{\hat{y}_{t,n-1}\geq\hat{y}_{t,n}}+(1-q)\cdot(\hat{y}_{t,n}-\hat{y}_{t,n-1})\cdot\mathbb{I}_{\hat{y}_{t,n-1}<\hat{y}_{t,n}}\right)\mbox{,}$ (4)
$\text{MQC}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QC}(q)\mbox{.}$ (5)
@@
The Quantile Change (QC) metric is a measure of the change in forecasted quantiles between two consecutive forecast origins. Given two forecast origins, $n$ and $n-1$, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous forecasts $\hat{y}_{t,n-1}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{y}_{t,n}\approx\hat{y}_{t,n-1}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of temporal robustness in its probabilistic output. By aggregating this across a range of quantiles $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability across the entire forecast distribution, providing a comprehensive metric for probabilistic stability. One drawback of the MQC is that it is based solely on a finite set of quantiles rather than the full predictive distribution. As a result, it captures changes only at selected points in the distribution (e.g., median, tails), potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Moreover, through Conformal Inference, it is possible to calculate as many quantiles as desired at no computational cost, making MQC a pragmatic and model-agnostic metric for capturing distributional instability across time.
@@
We computed 13 quantiles: the median and 6 central prediction intervals (60%, 70%, 80%, 90%, 95% and 99%). Central intervals describe the forecast center, while wider intervals capture tail risks, essential for safety stock decisions in retail (Barrow & Kourentzes, 2016). To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used.
@@
We normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining windows. To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
@@ 3.4.2 Performance metrics
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (Makridakis et al., 2022b). Lower values indicate better performance.
To assess probabilistic accuracy, we evaluated the quantiles produced through the Conformal Inference framework using the Quantile Loss (QL) and Multi-Quantile Loss (MQL).
$\text{QL}=\frac{1}{h}\sum_{t=n+1}^{n+h}\left(q\cdot(y_{t}-\hat{y}_{t})\cdot\mathbb{I}_{y_{t}\geq\hat{y}_{t}}+(1-q)\cdot(\hat{y}_{t}-y_{t})\cdot\mathbb{I}_{y_{t}<\hat{y}_{t}}\right)\mbox{,}$ (7)
$\text{MQL}=\frac{1}{\mathcal{Q}}\sum_{q\in\mathcal{Q}}\text{QL}(q)\mbox{.}$ (8)
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the MQL was the official metric of the M5 Uncertainty competition (Makridakis et al., 2022c).
@@ 4 Results and discussion
This section discusses the empirical findings of our study, highlighting the interplay between accuracy, probabilistic performance, and stability cost across different retraining strategies and ensemble configurations. We extract insights from both the M5 and VN1 datasets to uncover consistent patterns as well as nuances unique to each dataset’s characteristics.
Table 3: Overall forecasting performance and stability for each method across datasets. RMSSE, MQL, sMAPC, and MQC. Minimum values per column are highlighted in bold.
@@
RMSSE MQL sMAPC MQC RMSSE MQL sMAPC MQC
LR 0.777 0.267 0.188 0.118 6.549 2.896 0.085 1.064
RF 1.868 2.590 0.092 1.322
XGBoost 0.755 0.258 0.075 0.115 1.890 2.469 0.146 1.101
LGBM 0.771 0.256 0.054 0.123 3.542 2.625 0.121 1.282
CatBoost 0.947 0.263 0.118 0.127 5.762 2.845 0.320 1.435
MLP 0.821 0.281 0.500 0.109 1.543 2.492 0.458 0.844
LSTM 1.913 2.843 0.558 1.020
TCN 0.865 0.290 0.499 0.114 1.913 2.843 0.600 1.020
NBEATSx 0.815 0.279 0.547 0.110 1.698 2.626 0.596 1.532
NHITS 0.828 0.284 0.566 0.112 1.699 2.632 0.605 1.592
Ens2A 0.757 0.255 0.059 0.117 1.472 2.369 0.503 1.079
Ens3A 0.757 0.256 0.089 0.117 1.517 2.410 0.514 1.200
Ens4A 0.758 0.249 0.095 0.113 1.524 2.386 0.167 1.180
Ens5A 0.763 0.251 0.099 0.111 1.544 2.375 0.164 1.153
@@
Table 3 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets. For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, MQL (Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and MQC (Multi-Quantile Change) to understand the stability in probabilistic terms. Overall 111 The overall results are based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1, which is commonly considered the standard in both theoretical literature and practical forecasting applications. , the models evaluated achieved better stability results on the M5 dataset than on VN1. Several factors may account for this difference, such as the larger dataset size and higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of model training. Moreover, in terms of point forecasting, machine learning models are much more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates. On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
In general, ensembling seems to be a good solution to produce stable predictions. However, combining the most accurate models is often not sufficient to reduce the instability. Indeed, ensembles show increasing returns in stability as more models are added, as opposed to the well-known phenomenon in the ensemble literature, where accuracy improvements show diminishing returns from adding more models (Zanotti, 2025a). Interestingly, larger ensembles are usually more effective for stability purposes, implying that model diversity may be a more relevant criterion to create a forecast combinations, in contrast to model performance. This pattern is particularly true for probabilistic forecasting, but it is also evident in point predictions. For instance, in the VN1 dataset, the Ens4A and Ens5A combinations achieve good stability results simply because they contain both deep learning and machine learning models, extending the combination’s diversity. Indeed, even if the deep learning models’ instability is on average 60%, it is enough to add one single machine learning model to obtain an improvement of almost 45%.
@@
Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1. As expected, the sMAPC profiles are non-increasing functions of the retraining frequencies. The stability remains practically the same across update frequencies or even improves as the retraining period increases. Indeed, moving from high to low retraining frequencies, the stability of most global models improves compared to the baseline. In particular, for some models (i.e. CatBoost, LSTM, and TCN) the instability reduction is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability.
@@
In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting. In this context, we observe that, for the M5 dataset, the stability (as measured by Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models. For the VN1 dataset, instead, we observe an almost convex relationship between the stability and the retraining frequency. On average, models’ stability improves for low retraining scenarios and then starts deteriorating around $r=10$, implying that it is possible to avoid retraining up to 10 weeks without negatively impacting on the forecast stability. For both datasets, these results are confirmed by the statistical tests on MQC over the different retrain periods (see supplementary materials).
@@
Figure 4: MQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
@@
Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models. Lower retraining frequencies preserve both accuracy and stability, while offering substantial savings in computational resources.
@@
Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for various ensemble configurations. The results on the M5 dataset indicate that reducing retraining frequency may negatively affect forecast stability. However, this impact is below 1% for most retraining scenarios. Indeed, statistical tests (in supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining. The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining frequency in terms of forecast stability. Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM that where both the most accurate and the most stable global forecasting models on that set of data. Nonetheless, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones, especially at lower retraining frequencies, suggesting that model diversity plays a critical role in stabilizing ensemble forecasts when. Figure 6 further supports these observations in the probabilistic domain. First of all, probabilistic forecasts of ensembles are generally more stable than that of the base models across retraining frequencies, especially in M5, where MQC variations are within 1 percentage point. Moreover, in the VN1 dataset, the almost convex relationship between stability and the retraining frequency is flattened, because the variations are more restrained as the retraining period increases, and the decline in stability happens only from $r=26$ onwards. Then, larger ensembles provide consistently more stable probabilistic performance across most retraining scenarios, confirming the role of the diversity composition of ensembles even in the probabilistic setting. Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or MQC, is that it greatly reduces the variations across the different retraining frequencies, reducing both positive and negative stability changes.

### v2

@@ 3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir & Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
@@
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency. A step size of one is also relevant in the context of stability evaluation, since multiple forecasts for the same target are required to compute stability metrics.
Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
M5 VN1
Frequency ($f$) daily (7) weekly (52)
Train size ($n$) $\geq 730$ $\geq 157$
Test size ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b). Each scenario $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on aligned datasets, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. Following Zanotti (2025b), we treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
@@ 3.4.1 Stability metrics
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and n1n-1 respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply less unstable predictions.
@@
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios 222 One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models. These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022). However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research. . Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic quantile forecasts, we proposed a new measure, the Multi-Quantile Change (MQC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
$\text{QC}(\alpha)=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big{(}\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big{)}\mbox{,}$ (4)
$\text{MQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{QC}(\alpha)\mbox{,}$ (5)
@@
where $\mathcal{Q}$ is the set of possible quantile levels $\alpha$. The Quantile Change (QC) metric is a measure of the change in forecasted quantiles, $\hat{q}_{t}^{(\alpha)}$, between two consecutive forecast origins. Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of stability in its probabilistic output. By aggregating this across a set of quantile levels $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability, providing a comprehensive metric for probabilistic stability.
One drawback of the MQC is that it is a scale-dependent measure of instability. To solve this issue, we also propose a scaled version of the MQC, the Scaled Multi-Quantile Change (SMQC), defined as
$\text{SQC}(\alpha)=\frac{\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big{(}\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big{)}}{\frac{1}{n-s}\sum_{t=s+1}^{n}\left|y_{t}-y_{t-s}\right|}\mbox{,}$ (6)
$\text{SMQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQC}(\alpha)\mbox{,}$ (7)
where we adopted the in-sample one-step-ahead seasonal naive forecast as a scaling factor, to enable the effective application of this metric to forecasting models by ensuring that the results are not biased by the scale of individual time series.
To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$). KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback & Leibler, 1951). However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles. Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it again relies on full distributions (Peyré & Cuturi, 2019). The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need of full predictive densities. Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Additionally, given that it is based on the pinball loss, the MQC is inherently asymmetric, reflecting the fact that quantile forecasts are directional. For example, an unstable 90% quantile forecast may have stronger implications on inventory or risk management than a similarly unstable 60% quantile forecast. This implies that MQC emphasizes directional changes when measuring stability, in contrast to symmetric distances (like $W_{2}$), which treat deviations identically regardless of the quantile’s position. Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions. Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
@@
We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, essential to determine the safety stock levels in retail demand forecasting (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution. To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used 333 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 30,490 to 28,298 for the M5 dataset. .
@@
We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods. To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
@@ 3.4.2 Performance metrics
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022b). Lower values indicate better performance.
To assess probabilistic accuracy, we evaluated the quantiles produced using a scaled version of the Multi-Quantile Loss. The Scaled Quantile Loss, also known as (Scaled) Pinball Loss, and the Scaled Multi-Quantile Loss are defined as:
$\text{SQL}(\alpha)=\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}\Big{(}\alpha\cdot(y_{t}-\hat{q}_{t}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}\leq y_{t}}+(1-\alpha)\cdot(\hat{q}_{t}^{(\alpha)}-y_{t})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}>y_{t}}\Big{)}}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (9)
$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQL}(\alpha)\mbox{,}$ (10)
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted and with $s=1$) (Makridakis et al., 2022c).
@@ 4 Results and discussion
This section discusses the empirical findings of our study, highlighting the interplay between accuracy and stability across different retraining strategies.
Table 2: Overall forecasting performance and stability for each method across datasets based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1. Minimum values per column are highlighted in bold.
@@
RMSSE SMQL sMAPC SMQC RMSSE SMQL sMAPC SMQC
LR 0.777 0.267 0.188 0.110 6.549 4.286 0.085 1.560
RF 1.868 1.220 0.092 0.729
XGBoost 0.755 0.245 0.075 0.100 1.890 1.503 0.146 0.629
LGBM 0.771 0.246 0.054 0.118 3.542 2.318 0.121 1.484
CatBoost 0.947 0.277 0.118 0.156 5.762 4.233 0.320 2.276
MLP 0.821 0.263 0.500 0.083 1.543 0.806 0.458 0.219
LSTM 1.913 0.892 0.558 0.242
TCN 0.865 0.270 0.499 0.083 1.913 0.891 0.600 0.241
NBEATSx 0.815 0.262 0.547 0.084 1.698 0.991 0.596 0.765
NHITS 0.828 0.266 0.566 0.087 1.699 0.992 0.605 0.755
Ens2A 0.757 0.243 0.059 0.108 1.472 0.826 0.503 0.441
Ens3A 0.757 0.248 0.089 0.107 1.517 0.867 0.514 0.539
Ens4A 0.758 0.240 0.095 0.099 1.524 0.926 0.167 0.545
Ens5A 0.763 0.241 0.099 0.095 1.544 1.013 0.164 0.540
@@
Table 2 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets. For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, SMQL (Scaled Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and SMQC (Scaled Multi-Quantile Change) to understand the stability in probabilistic terms. Overall, the models evaluated achieved better stability results on the M5 dataset than on VN1. Several factors may account for this difference, such as the larger dataset size and a higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of the analysis. Moreover, in terms of point forecasting, machine learning models are more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates. On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
In general, ensembling seems to be a good solution to produce stable predictions when compared to base models. However, larger ensembles do not necessarily imply more stability, and it is not clear whether model diversity may have an impact on ensembles stability of predictions.
@@
Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1. As expected, the sMAPC profiles are non-increasing functions of the retraining scenario. The stability remains practically the same across update frequencies or even improves as the retraining period increases. Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline 444 Some models, like LR and MLP, show slightly decreasing stability (higher sMAPC) with an increasing retraining period. This can be due to many factors (e.g. sensitivity to small data changes, lack of robustness to stale parameters), but further investigation is necessary. . In particular, for some models (i.e. CatBoost, LSTM, and TCN), the instability reduction is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability.
@@
In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting. In this context, we observe that, for the M5 dataset, the stability (as measured by Scaled Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models 555 There are a few strange spikes in the SMQC profiles of some models. These can be due to different factors (e.g. inappropriate hyperparameter values), but further investigation is necessary. . For the VN1 dataset, instead, we observe some differences among the models. Some of them show constant improvements in stability as the retraining scenario increases (i.e. NBEATS, NHITS, and CatBoost), while most of them do not exhibit any significant enhancements or deteriorations in the retraining intervals. For both datasets, these results are confirmed by the statistical tests on SMQC over the different retraining periods (see Supplementary materials).
@@
Figure 4: SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
@@
Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models. Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources.
@@
Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for the various ensemble configurations. The results on the M5 dataset indicate that reducing retraining scenario may negatively affect forecast stability. However, this impact is below 1% for most retraining scenarios. Indeed, statistical tests (in Supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining. The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining scenario in terms of forecast stability. Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM, that where both the most accurate and the most stable global forecasting models on that set of data. On the contrary, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones in the VN1 dataset, suggesting that when base models are less accurate, bigger ensembles may be necessary to improve forecast stability.
Figure 6 further supports these observations in the probabilistic domain. First of all, probabilistic forecasts of ensembles are generally more stable than those of the base models across retraining frequencies, especially in M5, where SMQC variations are within 1 percentage point. In the VN1 dataset, the stability improvement is evident as the models are retrained less frequently. Then, from a probabilistic perspective, all the different ensembles perform similarly in terms of stability of the predictions, without showing any relevant benefit of having larger ensembles with respect to smaller ones. Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or SMQC, is that it greatly reduces the variations across the different retraining scenarios, reducing both positive and negative stability changes.